### v2

# Code Digital Twin: Empowering LLMs with Tacit Knowledge for Complex Software Maintenance

Abstract.
While large language models (LLMs) have demonstrated promise in software engineering tasks like code completion and generation, their support for the maintenance of complex software systems remains limited. These models often struggle with understanding the tacit knowledge embedded in systems, such as responsibility allocation and collaboration across different modules. To address this gap, we introduce the concept and framework of Code Digital Twin, a conceptual representation of tacit knowledge that captures the concepts, functionalities, and design rationales behind code elements, co-evolving with the software. A code digital twin is constructed using a methodology that combines knowledge extraction from both structured and unstructured sources—such as source code, documentation, and change histories—leveraging LLMs, static analysis tools, and human expertise. This framework can empower LLMs for software maintenance tasks such as issue localization and repository-level code generation by providing tacit knowledge as contexts. Based on the proposed methodology, we explore the key challenges and opportunities involved in the continuous construction and refinement of code digital twin.
copyright: none
1. Introduction
Recent advancements in large language models (LLMs) have demonstrated their impressive capabilities in software engineering (SE) tasks, such as code completion, bug detection, and software documentation. These models excel at understanding and generating code, streamlining development cycles, and assisting developers with routine tasks. Leveraging vast amounts of public domain data, LLMs have notably improved productivity and code quality in simpler, well-defined software systems. However, amidst rapid AI advancements dominated by LLMs, claims about “generative AI ending programming” and “revolutionizing software paradigms” abound. This hype may mislead enterprise leaders into expecting “10x efficiency gains”, leading to unrealistic demands on development teams. Consequently, investments flood into LLM fine-tuning and tool procurement, while essential long-term initiatives, such as building robust CI/CD pipelines, digital foundations, and knowledge repositories, are neglected. Current research (including industry efforts) excessively prioritizes the “glamour effect” of LLMs and AI agents, paying scant attention to software engineering’s intrinsic challenges.
In enterprise software development, tasks predominantly revolve around maintenance. Even “adding new code” is essentially a maintenance task: new code is rarely written on a blank canvas; instead, it must be integrated into proper positions within an existing software architecture through targeted additions and modifications. This process demands rigorous change impact analysis and regression testing to mitigate potential side effects. As software systems increase in complexity, LLMs face significant challenges. While they are effective in tackling individual tasks, their ability to support the maintenance and evolution of large, intricate systems is limited. LLMs struggle to capture the tacit knowledge essential to understanding complex software systems—such as design rationales, high-level functionalities, and system interdependencies—which is often undocumented or informally communicated. This challenge is exacerbated by knowledge decay, where critical context gradually fades or becomes fragmented across developer memories, scattered discussions, and incomplete documentation. As a result, LLMs lack the holistic, system-wide perspective needed to manage the intricate relationships and ongoing changes inherent in software evolution.
[…]
To address these challenges, we propose the concept and framework of Code Digital Twin, a conceptual representation of tacit knowledge that encapsulates the concepts, functionalities, and design rationales behind code elements, co-evolving with the software. Code Digital Twin addresses knowledge decay by extracting, accumulating, and continuously updating high-level knowledge about code through bidirectional mapping and co-evolution with the codebase. The framework structures tacit knowledge into two key categories: the artifact-oriented backbone, which establishes a foundational structure by linking concepts and functionalities to software artifacts, and the rationale-centric explanations, which enrich this foundation by capturing the decisions that have shaped the system’s design and evolution. We elevate software knowledge accumulation to a higher purpose: codifying knowledge about such ultra-complex systems becomes part of humanity’s civilizational record.
The methodology for realizing a Code Digital Twin combines structured and unstructured knowledge representations, integrating knowledge graphs, frames, and textual descriptions to capture domain concepts, functionalities, and rationales. The construction pipeline consists of three key phases: artifact-oriented backbone knowledge extraction, which employs top-down schema-based and schema-free extraction from documentation and bottom-up functionality summarization from source code; rationale-centric explanatory knowledge extraction, where LLMs extract and structure design rationales from unstructured sources like mailing lists and discussions using predefined knowledge frame schemas; and artifact-knowledge reflection construction, which links extracted knowledge across software artifacts and elements to create a cohesive representation. This pipeline supports incremental updates, ensuring that the Code Digital Twin evolves in sync with software changes, maintaining continuous alignment with the latest development state.
By providing integrated knowledge and ensuring continuous alignment with the evolving software, the Code Digital Twin enhances collaboration between LLMs and developers. It empowers LLMs to assist in software maintenance through various modes, including software question answering, context-aware development co-pilot, and autonomous development auto-pilot. This framework presents an innovative approach for the systematic accumulation of tacit knowledge in long-term software evolution, enabling more efficient maintenance and adaptation over time.
2. Motivation
Software systems are growing increasingly complex, while the pool of developers continues to evolve and fluctuate. This dynamic landscape makes software development more challenging, with essential knowledge becoming harder to retain and apply. A notable example is the Linux kernel, which remains under rapid development (Linux, 2025), with over 243,000 contributors producing more than 1,060,172 commits in 2021 alone. Its global, collaborative ecosystem introduces significant complexity, compounded by the continual influx and departure of developers. Software maintenance, a vital yet difficult phase of the development lifecycle, involves tasks such as diagnosing issues, assessing their impact, and integrating new code into existing architectures. These tasks require a deep understanding of both the system’s internal logic and external operating contexts. Despite recent advancements, evaluations such as those from OpenAI (Samuel et al., 2025) indicate that state-of-the-art LLMs remain limited in their ability to address complex, real-world maintenance challenges effectively.
2.1. Challenges in Complex Maintenance Tasks
Tacit Knowledge is Difficult to Capture. A significant portion of essential software knowledge is tacit—residing not in documentation or code, but in the minds of developers or in informal artifacts like whiteboards and meeting discussions. While LLMs excel at leveraging publicly available corpora, they falter when such tacit knowledge is missing or poorly documented. Critical insights into design rationale, historical decisions, and architectural intent are often unavailable for machine learning systems to absorb. This knowledge is also domain-specific, continuously evolving, and often requires human interpretation—making it unsuitable for passive ingestion via fine-tuning alone.
Internal and External Complexity. Software complexity manifests in both internal structures and external dependencies. Internally, systems are built upon intricate design decisions that result in deep interdependencies across the codebase. Externally, they must operate within a broader ecosystem—hardware, networks, runtime environments, and user scenarios—all of which influence design and maintenance. Contrary to simplified metaphors like “pattern soup”, software is less like a collection of reusable motifs and more akin to a precision instrument: highly engineered, tightly coupled, and often opaque. Its complexity exceeds that of mechanical systems due to four defining traits:
  • Inherent Ambiguity: Software encodes logic and intent, where small errors can have cascading effects, and boundaries between components are often ill-defined.
  • Contextual Dependencies: Code interacts with systems and conditions beyond itself, making isolated reasoning insufficient.
  • Internal Entanglement: High-level requirements or bug reports may map ambiguously to deeply entangled code, complicating traceability.
  • Evolutionary Uniqueness: Software evolves through decentralized contributions, leading to technical debt, duplication, and legacy structures.
LLMs lack the high-level, abstract reasoning and system-level perspective necessary to navigate such intricacies. They are not equipped to internalize the broader context in which systems operate or the creative processes that drive architectural decisions.
The “LLM Catch-All” Fallacy. A prevalent misconception in the field is the belief that LLMs can serve as a universal solution for software engineering by simply ingesting all relevant knowledge through pretraining or fine-tuning. This so-called LLM catch-all view treats software as a sequence of patterns, suggesting that intelligent development is merely a matter of recognizing and reproducing them. It implies that “all code is derivative”, downplaying the originality, structure, and judgment involved in real-world software work. However, this approach has fundamental flaws:
  • Flattened Information: Code without contextual metadata lacks essential abstractions and decision-making traces.
  • Absence of Design Intelligence: LLMs cannot infer the creative, system-wide reasoning behind design decisions from isolated tokens.
To enable intelligent support for tasks like root cause analysis, change planning, or impact estimation, we must move beyond superficial pattern recognition. These tasks require System 2 thinking—deliberate, structured, and cognitively intensive reasoning—something current LLMs are not equipped to perform independently.
2.2. Toward a New Methodology: Code Digital Twin
Given these challenges, LLMs are insufficient to support the development and maintenance of complex software systems on their own. Hence, the LLM catch-all approach fails for complex software. A new methodology is needed to systematically capture, organize, and maintain the tacit knowledge that underpins software evolution. The Code Digital Twin framework addresses this need by providing a structured representation of software knowledge that co-evolves with the system. Success hinges on:
  • Continuously codifying knowledge from development and evolution.
  • Structuring knowledge (e.g., via graphs).
  • Enabling visible, collaborative curation by developers.
Code Digital Twin is knowledge engineering for ultra-complex systems. But why “digital twin” for already-digital software? Code is the physical layer (files, classes, functions), while the Digital Twin captures the logical layer (requirements, design, decisions)—mapping intent to implementation. For cloud-native systems, we extend this to System Digital Twin, encompassing deployment topology, runtime environments, configuration states, and live updates. For monolithic systems (e.g., Linux), Code Digital Twin suffices—focusing on high-level abstractions and evolution. By continuously codifying knowledge, structuring it via graphs and frames, and enabling collaborative curation by developers, the Code Digital Twin empowers LLMs to assist in complex maintenance tasks more effectively, ensuring that critical knowledge is preserved and accessible for the long-term sustainability of software systems.
3. Code Digital Twin Framework
A code digital twin is a conceptual representation of tacit knowledge that encapsulates the concepts, functionalities, and design rationales behind code elements, co-evolving with the software. It enables efficient and effective software maintenance by providing on-demand, integrated insights at various granularities. Figure 1 provides an overview of the code digital twin’s construction process and its application paradigm.
Figure 1. The Overview of Code Digital Twin Framework
3.1. Definition
Key terminologies in the code digital twin framework are defined as follows:
Software Artifacts and Related Assets. These are the “physical” resources of a software system, directly accessible by developers, including source code, data files, commit histories, issue tracks, change logs, etc.
Key Knowledge Elements. These are high-level, conceptual assets underlying software, including domain concepts, functionalities, and rationales (design decisions at various levels).
  • Domain Concepts are background concepts for understanding the requirements and designing the solutions of a software system (e.g., basic OS concepts for Linux kernel).
  • Functionalities describe core functionalities implemented in the source code and documentation, often spanning multiple functions or packages.
  • Rationales explain design decisions, such as API choices or module partitioning, and are typically buried in unstructured assets like commit messages and issue threads.
Code Digital Twin. This model integrates key knowledge elements and their relationships through a unified representation closely linked to software artifacts. It consists of the following components:
  • Artifact-Oriented Backbone. A high-level representation oriented to software artifacts using concepts and functionalities, providing a shared, basic understanding of the system.
  • Rationale-Centric Explanations. Detailed explanatory rationales linked to the artifact-oriented backbone knowledge, supporting in-depth understanding of the system and decision-making.
  • Artifact-Knowledge Reflection and Co-Evolution. The backbone knowledge and explanatory rationales evolve alongside the software artifacts, ensuring continuous alignment.
3.2. Methodology
We propose a methodology for realizing the code digital twin, including principles and guidelines from a conceptualization and knowledgefication perspective.
3.2.1. Knowledge Representation.
The code digital twin leverages various knowledge representation methods to support the diverse categories of knowledge elements:
  • Structured Representation. Structured representations, such as knowledge graphs, knowledge cards, and frames, model relationships and hierarchies within knowledge units, providing a formal schema for organizing and querying the knowledge. For example, we can define the schema for domain concepts and their relationships, enriched with domain expertise (e.g., categorizing domain concepts and relationships, defining top-tier concepts, based on empirical guidelines such as open coding). Additionally, knowledge frames can be defined with necessary properties to store functionalities and different types of rationales. We can also integrate interaction schemas between knowledge graphs and frames by defining unified identities for knowledge units (e.g., concepts and functionalities).
  • Unstructured Representation. Textual descriptions, including natural language explanations and unstructured documents, offer flexible, context-rich representations that are not bound by predefined schemas, enabling more nuanced and descriptive knowledge capture. For instance, original functionality descriptions can be extracted from feature lists, and explanatory sentences for domain concepts can be stored.
  • Synergistic Representation. This approach integrates both structured and unstructured knowledge representations to create a comprehensive view, combining the strengths of both methods to enhance overall understanding and utility of the knowledge.
3.2.2. Construction Pipeline.
The code digital twin is constructed through a synergistic pipeline that leverages large language models (LLMs) and static program analysis.
  • Artifact-Oriented Backbone Knowledge Extraction: This phase consists of two parts: top-down extraction of concepts and functionalities from documentation with the assistance of human expertise, and bottom-up functionality summarization from source code. In the top-down extraction, schema-based methods using LLMs are employed to extract concepts and their predefined relationships from documentation texts and to structure the original functionality descriptions into predefined frame formats. This is followed by schema-free extraction with LLMs to capture additional backbone knowledge, such as open relationships between concepts. In the bottom-up extraction, LLMs are used to hierarchically summarize functionalities from functions, function call chains, modules and other cross-cutting code elements.
  • Rationale-Centric Explanatory Knowledge Extraction: This phase uses LLMs to extract instances of rationales based on predefined knowledge frame schemas. Initially, LLMs identify relevant information from a large, unstructured corpus (e.g., extensive mailing discussions). This information is then structured into predefined frame formats by using LLMs.
  • Artifact-Knowledge Reflection Construction: This phase utilizes LLMs to link concepts across artifacts and knowledge elements and to map functionalities between software artifacts and knowledge elements.
3.2.3. Co-Evolution and Incremental Update
The pipeline must support incremental updates to maintain alignment with the real-time status of the software artifacts and related assets. As the software evolves, the knowledge and representation within the code digital model are updated synchronously.
3.3. Application and Impact
The code digital twin enables intelligent applications for complex software systems, categorized into three modes of collaboration between developers, LLMs, and the digital twin:
  • On-Demand Knowledge Visualization. This application provides dynamic, context-sensitive visualizations of software knowledge, offering developers instant insights into the system’s structure, behavior, and interdependencies, drawn from multiple sources, to aid in analysis and decision-making.
  • Context-Aware Development Co-Pilot. The digital twin acts as an intelligent knowledge backbone for an AI assistant, providing task-specific insights and knowledge snippets to help LLMs navigate challenges and complete tasks more accurately.
  • Autonomous Development Auto-Pilot. In this mode, LLM-based agents autonomously monitor the software lifecycle using the code digital twin, detecting and resolving issues to ensure stability and reduce the need for manual oversight.
Overall, the code digital twin will revolutionize the management and evolution of long-term complex software systems, enabling more efficient development, maintenance, and adaptation over time.
4. Research Agenda
To implement the code digital twin framework, we outline the following research agenda and identify actionable challenges and opportunities.
4.1. Synergistic Knowledge Representation
The goal is to develop a scalable framework for representing complex software knowledge in a digital twin model. This framework should integrate structured representations to organize code, design decisions, documentation, and historical changes, alongside unstructured representations for richer context. It should emphasize higher-level relationships and adapt to software changes. The key challenges and opportunities are outlined below:
  • Structured Knowledge Representation and Integration. Developing a flexible, scalable knowledge model for complex software requires integrating diverse knowledge sources, such as domain-specific graphs and ontologies, to map relationships between components and decisions. The challenge is to align disparate sources, ensuring consistency and improving the digital twin’s accuracy. Research should focus on developing knowledge fusion techniques to merge these sources efficiently.
  • Integrating Structured and Unstructured Knowledge. Combining structured data (e.g., code, class hierarchies) with unstructured data (e.g., documentation, design rationale) into a unified model is complex. Actionable research should focus on creating hybrid models that integrate knowledge graphs with NLP-driven text extraction, enabling both automated LLM reasoning and human interpretation for software development and maintenance.
  • Managing Ambiguity and Knowledge Versioning. Handling uncertain or evolving knowledge, such as design decision trade-offs, is a major challenge. Research should focus on using techniques like probabilistic reasoning and knowledge versioning to address ambiguity, ensuring the digital twin remains adaptable and coherent as the software evolves.
4.2. Automatic Construction Pipeline
The goal is to establish an efficient, accurate, and cost-effective pipeline for automatically extracting and structuring knowledge from software artifacts. This pipeline should integrate with the digital twin, ensuring continuous updates as the software system evolves. The key challenges and actionable opportunities are outlined below:
  • Ensuring Accurate Knowledge Extraction. Extracting relevant, high-quality knowledge from diverse sources while minimizing noise is a critical challenge. To improve accuracy, focus on refining LLM-based extraction through targeted prompt engineering and fine-tuning for specific domains. This can improve the processing of both structured and unstructured data, such as source code, feature descriptions, and mailing list.
  • Scalability and Efficiency of the Pipeline. Scaling the extraction pipeline for large, complex software systems while maintaining accuracy and speed is challenging. Focus on implementing distributed processing to enable parallel extraction and transformation of large data sets. Additionally, leverage incremental learning and batch processing to ensure scalability without compromising extraction quality over time.
  • Adapting to Evolving Software Artifacts. Ensuring the extraction pipeline adapts to frequent software updates, including changes to code and documentation, is a key challenge. Implement real-time synchronization techniques to capture and integrate these changes into the digital twin as they occur, maintaining accurate and up-to-date knowledge representation.
4.3. Human-in-the-Loop Knowledge Accumulation
The goal is to leverage human input, through crowdsourcing and expertise, to enhance and update the digital twin. This approach fills gaps, ensures completeness, and creates a feedback loop where human insights complement automated knowledge extraction. Integrating human expertise makes the digital twin more dynamic, accurate, and reflective of real-world systems. The key challenges and opportunities are outlined below:
  • Encouraging and Ensuring Quality Human Contributions. Motivating developers and experts to contribute is challenging without clear incentives. Research should focus on developing effective incentive mechanisms, such as reputation systems or gamification, to drive participation. Additionally, ensuring contribution quality requires systematic evaluation, such as using metrics (e.g., bug reduction or performance improvements), peer reviews, and expert validation to maintain accuracy and reliability.
  • Designing User-Friendly Interfaces for Contributors. Creating intuitive, seamless interfaces for contributors is critical. Solutions like IDE plugins that unobtrusively capture programming activities (e.g., code maintenance patterns) will enable passive knowledge accumulation. Interfaces should offer guidance, progress tracking, and contextual feedback, focusing contributions on high-impact areas while remaining accessible to various types of participation.
  • Maintaining Knowledge Completeness Over Time. As software systems evolve, keeping the accumulated knowledge up-to-date is challenging. Research should focus on developing automated tools for real-time synchronization with evolving software systems and establishing periodic reviews to identify and address knowledge gaps, ensuring the digital twin stays relevant and current.
4.4. Integration with Large Language Models
The goal of integrating the code digital twin with LLMs is to improve software development through intelligent collaboration, problem-solving, and predictive analysis. This integration combines the deep context in the digital twin with the language processing capabilities of LLMs to enhance software maintenance, issue resolution, and decision-making. The challenges and opportunities are outlined below:
  • Software Question Answering. A key challenge is dynamically synthesizing contextual software views, such as architectural impact maps and rationale traces, that meet developers’ evolving needs. To address this, research should focus on developing hybrid models that combine LLMs’ natural language understanding with the digital twin’s knowledge bases, enabling interactive visualizations. These models should translate natural language queries into actionable visual narratives, highlighting hidden relationships while managing information overload through targeted abstraction.
  • Context-Aware Development Co-pilot. Balancing developer autonomy with LLM-driven suggestions is challenging, especially in modeling the semantics of code (from the digital twin) and the developer’s intentions (e.g., design exploration or issue localization). A promising solution is to develop semantic retrieval techniques that extract relevant knowledge snippets from the code digital twin to support task completion.
  • Autonomous Development Auto-pilot. Enabling LLMs to autonomously manage software tasks (e.g., bug fixing, refactoring) while adhering to project goals and design principles is a key challenge. Unchecked autonomy could lead to violations of system invariants or inefficient solutions. To address this, research should focus on integrating development tools (e.g., static analyzers, linters) with the code digital twin, ensuring real-time validation and contextual guidance for LLMs or LLM-based agents to support autonomous decision-making and maintain alignment with project objectives.
4.5. Co-Evolution of Code Digital Twin
The goal of a co-evolving twin is to ensure that the code digital twin remains an up-to-date, relevant, and dynamic representation of the evolving software system. This involves continuous synchronization with the codebase, incorporating developer feedback, adapting to system changes, and maintaining high-quality and consistent knowledge. Below are the key challenges and opportunities:
  • Ensuring Continuous Synchronization and Adaptation. A primary challenge is keeping the code digital twin in sync with the continuously evolving codebase, which includes frequent updates, new features, and design changes. This can be achieved by automating synchronization through tools like continuous integration (CI) pipelines, change detection mechanisms, and incorporating developer feedback from issue trackers and version histories. Additionally, adopting modular frameworks and integrating automatic documentation can ensure the code digital twin dynamically adapts to system changes, maintaining relevance and accuracy over time.
  • Maintaining Knowledge Quality and Consistency. As the system evolves, outdated or inconsistent knowledge may accumulate in the code digital twin. Implementing automated checks, version tracking, and semantic validation can help ensure that the digital twin’s knowledge remains accurate and consistent over time, despite ongoing changes in the software system.
  • Tracking the Impact of Code Changes. Assessing the impact of software changes on the accumulated knowledge in the code digital twin can be challenging. Mechanisms that trace code changes and evaluate their effect on the digital twin’s knowledge will help ensure the digital twin remains responsive and accurate in reflecting the evolving system requirements.
5. Related Work
Repository-level code generation with LLMs. Repository-level code completion involves generating code based on cross-file context within a repository. Two main techniques assist LLMs in this task: Retrieval-Augmented Generation (RAG) (Wu et al., 2024; Liu et al., 2024a; Zhang et al., 2023; Liang et al., 2024), which retrieves similar code snippets, and static analysis (Agrawal et al., 2023; Wang et al., 2024; Liu et al., 2024a; Liang et al., 2024; Pei et al., 2023), which uses imported files or API lists. However, both approaches face challenges: retrieved snippets may not always help, and static analysis can be inaccurate, especially for dynamically typed languages. Despite advances in LLMs, accurate repository-level code completion remains difficult due to cross-file context complexities and limitations of current techniques.
Repository-level issue resolution with LLMs. Repository-level issue resolution is tackled through two main approaches: Agent-based (Yang et al., 2024; Zhang et al., 2024; Chen et al., 2024) and Pattern-based (Xia et al., 2024) methods. Agent-based approaches, like SWE-Agent (Yang et al., 2024), treat LLMs as agents to edit code, navigate repositories, and collaborate on issues. Pattern-based methods, such as AgentLess (Xia et al., 2024), involve retrieving relevant code context in the Search/Localize stage, followed by code modification in the Edit stage. Other works (Liu et al., 2024b; Ma et al., 2024; Ouyang et al., 2024) mine structural knowledge from code graphs to improve search results. However, the large scale of repositories and complex issues make it challenging for LLMs to capture implicit knowledge like long call chains or design patterns.
Repository comprehension with LLMs. Many researchers are leveraging the comprehension capabilities of LLMs alongside knowledge representation techniques to address repository comprehension tasks. For instance, MutableAI (MutableAI, 2025) enhances comprehension by converting codebases into Wikipedia-style articles with citations, enabling AI-powered experts to generate context-aware answers to repository queries. Similarly, CODEXGRAPH (Liu et al., 2024c) integrates LLMs with graph database interfaces derived from code repositories, allowing LLMs to construct and execute queries for efficient comprehension. These methods rely on the effectiveness of code wikis or graphs, highlighting the need for further research on how to better represent and use the information within code repositories to unlock the potential of LLMs.
Knowledge graphs for software engineering. Researchers have developed various knowledge graphs to aid software development tasks. API KGs (Li et al., 2018; Liu et al., 2019, 2020a; Ren et al., 2020; Peng et al., 2018; Liu et al., 2020b) summarize API knowledge, supporting tasks like migration and misuse detection. Software development concept KGs (Zhao et al., 2017; Wang et al., 2019, 2023b, 2023a; Xing et al., 2021; Liu et al., 2023) create domain glossaries. Programming task KGs (Liu et al., 2022; Sun et al., 2019) extract task attributes and relationships from tutorials, helping developers solve programming questions. Bug KGs (Xie et al., 2019; Su et al., 2021; Du et al., 2023) capture knowledge on bug causes and solutions. However, no knowledge base specifically summarizes design-related knowledge for long-term software maintenance tasks.
6. Conclusion
In this work, we introduce the Code Digital Twin, a conceptual framework designed to bridge the gap between LLMs and the tacit knowledge essential for complex software maintenance. Our proposed methodology integrates LLMs, static analysis, and human expertise to construct and continuously evolve this representation in alignment with software development. Through this framework, we aim to enhance the effectiveness of LLMs in understanding system-wide dependencies, design rationales, and evolving software artifacts. Looking ahead, we identify key challenges in ensuring the scalability, adaptability, and long-term evolution of the code digital twin, opening avenues for further research in knowledge-centric software maintenance.